Grade 12th passPhysical Chemistry

Calculate equilent concentration of magnesium chloride if 500ml solution contains 0.95gm Mgcl2​

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Pawan Prajapati
5 Years ago
Answer:

Explanation:

Molar mass of MgCl2 =95g/mol

No.of mol = 95g/95g/mol= 1mol

Molarity =( 1/500)×(1000) =2M

So

Equivalent concentration = n- factor × 2 N= 2×2 =4N

Thanks